2. Find differences and similarities of the following functions (1) y =x; (2) y=Vr?, y = Siri (3) y =]x], (4) and (5)
Answer #1

x 16-8)2 = y = 572 = 1x) (o if y = 2x2 = x 118-72 =) 7-8 = 8-7 - 14 harelato 16 a contradiction 3 y = lul = x=2 42 y = risu=